Factoring the quadratic: I believe that there is an error in the solution.
sin^2 theta + 4 sin theta + 3 = 0
I have tried bringing the 3 over to the right side - produces solution that is off the chart.
The solution provides sin theta = +1, and +3. However the work does not check out as you break the factoring apart and factor both terms separately.
Please advise...

Answer by rothauserc(4718)

sin^2 theta + 4 sin theta + 3 = 0
this can be factored into
(sin(theta) + 1)(sin(theta) + 3) = 0
therefore,
sin(theta) + 1 = 0
sin(theta) + 3 = 0